# Uniqueness Analysis of Controllability Scores and Their Application to Brain Networks

### v3

@@ II-A VCS and AECS
Note that $p\in X_{T}$ means that system (2) is controllable. Moreover, the constraint $p\in\Delta$ ensures that $p_{1},\ldots,p_{n}$ reflect the relative importance of the corresponding state nodes $x_{1},\ldots,x_{n}$.
For an optimal solution $p$ to Problem (9), we define the controllability score for state node $x_{i}$ as follows: if $h_{T}(p)=f_{T}(p)$, then $p_{i}$ is termed the volumetric controllability score (VCS); if $h_{T}(p)=g_{T}(p)$, then $p_{i}$ is the average energy controllability score (AECS). As explained in [19], VCS of each state node indicates its importance in enlarging the controllability ellipsoid
@@
AECS of each state node indicates its importance in steering the overall state to a point on the unit sphere. Hereafter, we refer to (9) as the finite-time controllability scoring problem (FTCSP).
Larger VCS and AECS values highlight its significant role in the system’s controllability. In fact, $W_{i}(T)\neq 0$ implies that for any $p\in X_{T}$,
@@
$(\nabla g_{T}(p))_{i}$ $=-{\rm tr}(W(p,T)^{-1}W_{i}(T)W(p,T)^{-1})<0,$ (13)
where $\nabla f_{T}$ and $\nabla g_{T}$ denote the gradients of $f_{T}$ and $g_{T}$, respectively. Because both gradients are strictly negative, increasing $p_{i}$ enhances the volume of the controllability ellipsoid $\mathcal{E}(p,T)$ and reduces the average energy required for state steering [19].
@@
Suppose that a solution to FTCSP (9) uniquely exists. If $\{p^{(k)}\}$ is a sequence generated by Algorithm 1 with $\varepsilon=0$, then
@@ II-B Existing results
The interpretation of Propositions 2, 3, 4 is as follows: Proposition 2 states that VCS and AECS can be uniquely defined for any asymptotically stable system (1) and any $T>0$. Proposition 3 suggests that VCS and AECS can be uniquely defined for many systems without imaginary-axis eigenvalues and for any $T>0$. Proposition 4 indicates that VCS and AECS can be uniquely defined for any Laplacian dynamics (1) with $A=-L$ from an undirected graph and for any $T>0$.
@@ II-C Limitations of existing results and objectives of this paper
    It remains unclear whether VCS and AECS can be uniquely defined for all linear systems of the form (1), particularly for Laplacian dynamics corresponding to directed graphs.
@@ III-B VCS and AECS in special cases
According to Theorem 2, if $A$ is close to a symmetric matrix, we can expect the VCS of each state node to be closer to ${\bf 1}/n$ compared to AECS.
@@
Corollary 1 states that for Laplacian dynamics on an undirected graph with $n$ nodes, the VCS of each state node is uniformly $1/n$.
@@ III-C Effect of the terminal time $T$
We investigate the effect of the terminal time $T$ on the optimal solution to FTCSP (9). That is, we resolve (iii), as mentioned in Section II-C. To this end, we define $Z_{i}(t):=\exp(At)e_{i}e_{i}^{\top}\exp(A^{\top}t)$. Then, the finite time Gramian $W_{i}(T)$ in (8) can be expressed as
@@
The finite time controllability Gramian $W(p,T)$ defined in (7) can be expressed as
@@
We have
$W_{i}(T)=Z_{i}(0)T+O(T^{2}),$ (31)
since (29) holds and
$Z_{i}(t)=Z_{i}(0)+\frac{\mathrm{d}Z_{i}}{\mathrm{d}t}(0)t+o(t),$
where $o(t)$ is a function that satisfies $\lim_{t\rightarrow 0}\frac{o(t)}{t}=0$. From (7), (31), and $Z_{i}(0)=e_{i}e_{i}^{\top}$, we obtain (30). $\Box$
III-C1 When $T$ is close to zero
@@
$\frac{\partial L_{f}}{\partial p_{i}}(p,\lambda)=-\frac{1}{p_{i}}+\lambda,\,\,\frac{\partial L_{g}}{\partial p_{i}}(p,\lambda)=-\frac{1}{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T}p_{i}^{2}}+\lambda$ (38)
for any $p\in\Delta_{\rm int}$ and $\lambda\in{\mathbb{R}}$. Thus, setting $\frac{\partial L_{f}}{\partial p_{i}}(p,\lambda)=0$ yields $\lambda=1/p_{i}$ for each $i=1,\ldots,n$, implying that the optimal solution satisfies $p_{1}=\cdots=p_{n}$. Combining this with the constraint $\sum_{i=1}^{n}p_{i}=1$, we obtain $p_{1}=\cdots=p_{n}=1/n$, which is an optimal solution to problem (34) with $\widetilde{h}_{T}(p)=\widetilde{f}_{T}(p)$. Similarly, setting $\frac{\partial L_{g}}{\partial p_{i}}(p,\lambda)=0$ yields $\lambda=1/({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T}p^{2}_{i})$ for each $i=1,\ldots,n$, implying that the optimal solution satisfies $p_{1}=\cdots=p_{n}$. Thus, $p_{1}=\cdots=p_{n}=1/n$, which is an optimal solution to problem (34) with $\widetilde{h}_{T}(p)=\widetilde{g}_{T}(p)$.
@@ III-C2 When $T$ is sufficiently large
When $T$ is sufficiently large, we must consider second term in (30), which renders the analysis considerably more challenging for a general matrix $A$ compared to the small $T$ case. Therefore, we assume that $A$ is stable. Here, we call $A$ stable if all eigenvalues of $A$ have negative real parts, and we say system (1) to be asymptotically stable if $A$ is stable.

### v4

@@ II-A VCS and AECS
Note that $p\in X_{T}$ means that the virtual system (2) is controllable. Moreover, the constraint $p\in\Delta$ ensures that $p_{1},\ldots,p_{n}$ reflect the relative importance of the corresponding state nodes $x_{1},\ldots,x_{n}$.
For an optimal solution $p$ to Problem (9), we define the controllability score for state node $x_{i}$ as follows: if $h_{T}(p)=f_{T}(p)$, then $p_{i}$ is termed the volumetric controllability score (VCS); if $h_{T}(p)=g_{T}(p)$, then $p_{i}$ is the average energy controllability score (AECS). As explained in [19], each VCS of each state node indicates its importance in enlarging the controllability ellipsoid
@@
Each AECS of each state node indicates its importance in steering the overall state to a point on the unit sphere. Hereafter, we refer to (9) as the finite-time controllability scoring problem (FTCSP).
Larger VCS and AECS values highlight their significant contribution to the system’s controllability. In fact, since $W_{i}(T)\neq O$ (because ${\rm tr}(W_{i}(T))>0$), for any $p\in X_{T}$, the $i$-th components of the gradients of $f_{T}$ and $g_{T}$ satisfy
@@
$(\nabla g_{T}(p))_{i}$ $=-{\rm tr}(W(p,T)^{-1}W_{i}(T)W(p,T)^{-1})<0.$ (13)
Because both gradients are strictly negative, increasing $p_{i}$ enhances the volume of the controllability ellipsoid $\mathcal{E}(p,T)$ and reduces the average energy required for state steering [19].
@@
Suppose that an optimal solution to FTCSP (9) is unique. If $\{p^{(k)}\}$ is a sequence generated by Algorithm 1 with $\varepsilon=0$, then
@@ II-B Existing results
The interpretations of Propositions 2, 3, 4 are as follows: Proposition 2 states that VCS and AECS are uniquely defined for any asymptotically stable system (1) and for all $T>0$. Proposition 3 suggests that VCS and AECS remain uniquely defined for any system whose eigenvalues lie off the imaginary axis, again for all $T>0$. Proposition 4 indicates that VCS and AECS are uniquely defined for any symmetric Laplacian dynamics (1) with $A=-L$ corresponding to an undirected graph, for all $T>0$.
@@ II-C Limitations of existing results and objectives of this paper
    It remains unclear whether VCS and AECS are uniquely defined for all linear systems of the form (1), especially in the case of non-symmetric Laplacian dynamics arising from directed graphs.
@@ III-B VCS and AECS in special cases
According to Theorem 2, if $A$ is close to being symmetric, we expect each node’s VCS to lie closer to the uniform value ${\bf 1}/n$ than its corresponding AECS value.
@@
Corollary 1 states that for symmetric Laplacian dynamics on an undirected graph with $n$ nodes, the VCS of each state node is uniformly $1/n$.
@@ III-C Effect of the terminal time $T$
We investigate the effect of the terminal time $T$ on the optimal solution to FTCSP (9). That is, we partially resolve (iii), as mentioned in Section II-C. To this end, we define $Z_{i}(t):=\exp(At)e_{i}e_{i}^{\top}\exp(A^{\top}t)$. Then, the finite time controllability Gramian $W_{i}(T)$ in (8) can be expressed as
@@ III-C1 When $T$ is close to zero
For the finite time controllability Gramian $W(p,T)$ in (7), there exists $\delta>0$ such that for all $T\in[0,\delta]$,
@@
Since $Z_{i}$ is analytic, the Taylor’s theorem implies
$Z_{i}(t)=Z_{i}(0)+\frac{\mathrm{d}Z_{i}}{\mathrm{d}t}(0)t+\int_{0}^{t}\frac{\mathrm{d}^{2}Z_{i}}{\mathrm{d}t^{2}}(\tau)(t-\tau)\mathrm{d}\tau.$
Since (29) holds,
$W_{i}(T)=$ $Z_{i}(0)T+\frac{1}{2}\frac{\mathrm{d}Z_{i}}{\mathrm{d}t}(0)T^{2}$
$+\int_{0}^{T}\int_{0}^{t}\frac{\mathrm{d}^{2}Z_{i}}{\mathrm{d}t^{2}}(\tau)(t-\tau)\mathrm{d}\tau\mathrm{d}t.$ (31)
Because the integrand $\frac{\mathrm{d}^{2}Z_{i}}{\mathrm{d}t^{2}}(\tau)(t-\tau)$ of the third term is integrable, the Fubini’s theorem guarantees that we can interchange the order of integration. Thus,
$\int_{0}^{T}\int_{0}^{t}\frac{\mathrm{d}^{2}Z_{i}}{\mathrm{d}t^{2}}(\tau)(t-\tau)\mathrm{d}\tau\mathrm{d}t$
$=$ $\int_{0}^{T}\frac{\mathrm{d}^{2}Z_{i}}{\mathrm{d}t^{2}}(\tau)\left(\int_{\tau}^{T}(t-\tau)\mathrm{d}t\right)\mathrm{d}\tau$
$=$ $\frac{1}{2}\int_{0}^{T}\frac{\mathrm{d}^{2}Z_{i}}{\mathrm{d}t^{2}}(\tau)(T-\tau)^{2}\mathrm{d}\tau.$ (32)
By substituting (32) into (31), we obtain
$\|W_{i}(T)-Z_{i}(0)T\|\leq$ $\frac{1}{2}\left\|\frac{\mathrm{d}Z_{i}}{\mathrm{d}t}(0)\right\|T^{2}$ (33)
$+\frac{1}{2}\int_{0}^{T}\left\|\frac{\mathrm{d}^{2}Z_{i}}{\mathrm{d}t^{2}}(\tau)\right\|(T-\tau)^{2}\mathrm{d}\tau.$
[…]
$\delta:=\min_{i\in\{1,\ldots,n\}}\delta_{i},\quad M:=\max_{i\in\{1,\ldots,n\}}\left\|\frac{\mathrm{d}^{2}Z_{i}}{\mathrm{d}t^{2}}(0)\right\|.$
Then, (33) implies that for any $T\in[0,\delta]$,
$\|W_{i}(T)-Z_{i}(0)T\|$ $\leq\frac{1}{2}\left\|\frac{\mathrm{d}Z_{i}}{\mathrm{d}t}(0)\right\|T^{2}+\frac{1}{3}MT^{3}$
$\leq CT^{2},$ (34)
where $C:=\frac{1}{2}\left\|\frac{\mathrm{d}Z_{i}}{\mathrm{d}t}(0)\right\|+\frac{1}{3}M\delta$. Because $Z_{i}(0)=e_{i}e_{i}^{\top}$ and $\|W(p,t)-T{\rm diag}(p_{1},\ldots,p_{n})\|\leq\sum_{i=1}^{n}|p_{i}|\|W_{i}(T)-Z_{i}(0)T\|$, (34) yields (30). $\Box$
@@
$\frac{\partial L_{f}}{\partial p_{i}}(p,\lambda)=-\frac{1}{p_{i}}+\lambda,\,\,\frac{\partial L_{g}}{\partial p_{i}}(p,\lambda)=-\frac{1}{Tp_{i}^{2}}+\lambda$ (41)
for any $p\in\Delta_{\rm int}$ and $\lambda\in{\mathbb{R}}$. Thus, setting $\frac{\partial L_{f}}{\partial p_{i}}(p,\lambda)=0$ yields $\lambda=1/p_{i}$ for each $i=1,\ldots,n$, implying that the optimal solution satisfies $p_{1}=\cdots=p_{n}$. Combining this with the constraint $\sum_{i=1}^{n}p_{i}=1$, we obtain $p_{1}=\cdots=p_{n}=1/n$, which is an optimal solution to problem (37) with $\widetilde{h}_{T}(p)=\widetilde{f}_{T}(p)$. Similarly, setting $\frac{\partial L_{g}}{\partial p_{i}}(p,\lambda)=0$ yields $\lambda=1/(Tp^{2}_{i})$ for each $i=1,\ldots,n$, implying that the optimal solution satisfies $p_{1}=\cdots=p_{n}$. Thus, $p_{1}=\cdots=p_{n}=1/n$, which is an optimal solution to problem (37) with $\widetilde{h}_{T}(p)=\widetilde{g}_{T}(p)$.
@@ III-C2 When $T$ is sufficiently large
Since Lemma 4 holds only for sufficiently small $T$, it cannot be applied when $T$ is large. As a result, the analysis for large $T$ is, in general, challenging.
Therefore, we assume that $A$ is stable. Here, we call $A$ stable if all eigenvalues of $A$ have negative real parts, and we say system (1) to be asymptotically stable if $A$ is stable.